OpenStudy (anonymous):

Help with supplementary geometry? When I use the formula, I end up with a decimal, did I do something wrong? Example: 8x+14 and 9x. so, 8x+14 + 9x= 180. When I combine like terms, I get decimals and end up getting the question wrong? 17x+14 = 180 -14 -14 = 17x = 166 When I divide it, I get 9.7? I really need help on this, it's really frustrating.

OpenStudy (anonymous):

What do you need help with?

OpenStudy (anonymous):

I edited my post.

OpenStudy (mertsj):

We need to see the original problem. Perhaps it is not the sum that is equal to 180.

OpenStudy (anonymous):

Well, that was an example, here is the problem.

OpenStudy (mertsj):

Yes. Add the two angles and the sum will be 180.

OpenStudy (anonymous):

I do this yet I get the wrong answer, I follow the procedure to do this problem. But, in the end, the answer doesn't reflect. The question to that problem is to solve for X.

OpenStudy (mertsj):

15x-20+35x=180 50x-20=180 50x=200 x=4

OpenStudy (anonymous):

Sorry, wrong equation, I just realized that

OpenStudy (anonymous):

here is the real problem.

OpenStudy (mertsj):

Alternate exterior angles are equal.

OpenStudy (anonymous):

I know that, just how do I solve for x though?

OpenStudy (mertsj):

Write the equation and post it. Then I will help you with it.

OpenStudy (anonymous):

The equation is: 8x + 14 + 9x = 180

OpenStudy (mertsj):

I thought we said that the two exterior angles are equal?

OpenStudy (mertsj):